OpenStudy (anonymous):

A ball has a radius of 18 cm. What is the approximate volume of the ball? Use 3.14 to approximate pi and express your answer in hundredths. cm3 @Hesan @MasterprojectHD I got 24416.64

OpenStudy (anonymous):

what did you do to solve your problem?

OpenStudy (anonymous):

you have 18cm and 3.14.

OpenStudy (anonymous):

would you times or divide, add or subtract?

OpenStudy (anonymous):

formula is 4/3pir^3

OpenStudy (kendricklamar2014):

A ball is the same thing as a sphere, so were going to find the volume of a sphere

OpenStudy (anonymous):

do you have answer choices? I just used the calculator with what you have just gave me and I got ((((0.0430020459

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I thought you had to go lol?!?!?!

OpenStudy (kendricklamar2014):

Formula:\[\left(\begin{matrix}4 \\ 3\end{matrix}\right) \pi r^3\]

OpenStudy (kendricklamar2014):

Now we substitute what we know:\[\left(\begin{matrix}4 \\ 3\end{matrix}\right) \pi 18^3\]

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}4 \\ 3\end{matrix}\right) \pi 18^3 = 24429.02\]

OpenStudy (anonymous):

5832

OpenStudy (anonymous):

Your correct.

OpenStudy (anonymous):

me or him?

OpenStudy (anonymous):

You

OpenStudy (kendricklamar2014):

Do you have any answer choices @cvliboi

OpenStudy (anonymous):

no

OpenStudy (anonymous):